# Quantum Optimization Benchmark Library The Intractable Decathlon

### v1

@@ 4.4.2 Problem statement
For $N\in\mathbb{N}_{+}$, we write $[N]=\{1,\ldots,N\}$. In the STPP, we are given an undirected simple graph $G=(V,E)$, and $N\in\mathbb{N}_{+}$ disjunct terminal sets $T_{1},\ldots,T_{N}$ with $T_{i}\subseteq V$ for all $i\in[N]$ and $T_{i}\cap T_{j}=\emptyset$ for $i,j\in[N]$, $i\neq j$. The goal is to find an edge set $S^{*}\subseteq E$ of minimal cardinality such that $(V(S^{*}),S^{*})$ is a forest, where $\bigcup_{i\in[N]}T_{i}\subseteq V(S^{*})$ and no vertex from $T_{i}$ is connected to any vertex of $T_{j}$ for all $i\neq j$. The general STPP with $N>1$, as defined above, requires the inclusion of several disjoint Steiner trees into the same graph. A solution is a forest, and each terminal set $T_{i}$ is connected by a single tree in the forest.
@@ 4.4.3 Instances
We generated 3-dimensional $S\times S\times L$ grid graphs with holes. All terminals are placed at the border of the top layer. We require all trees to be node-disjoint. The parameters for each problem are the size of the grid graph $S$, the number of layers $L$, the maximum number of terminals per net $T$, and the number of holes in the grid $H$. The holes can extend to multiple layers. An example of a $20\times 20\times 2$ instance with the corresponding optimal solution is given in Figure 7. In practice, the problem can also be defined with edge weights. For example, changing from one layer to another could be more expensive. However, since this does not make the problem more difficult in general, we kept it simple and set all edge weights to one.
Figure 8: (Steiner) Total runtime and gap to the best-known (or even provably optimal) solution, solved with Gurobi for different size $20$ instances with a timeout of 72000s (left) and different size $30$ instances with a timeout of 86400s (right). The ticked line represents the timeout, the crosses indicate the runtime, and the stars show the gap.
@@ 4.4.4 Classical Baseline
Using the multicommodity flow ILP formulation for the node-disjoint STPP given in [103], we ran Gurobi 11.0.2 on an AMD EPYC 7513 32-core processor with 64 threads to attempt to solve the smallest instances in our benchmark set (grid size 20 or 30). We used the model “as is”, without additional valid inequalities or decomposition strategies (the model is also provided in the repository [193]). Hardware information and results are summarized in Figure 8, whereas solution values and bound are given in Table 4. With this ILP model, we could solve most of the size 20 instances (with an average solution time of 53 minutes) within 20 hours, but only two of the size 30 instances within 24 hours (the solution times were 194 minutes and 624 minutes). Reimplementing the techniques of [103] using current solver technology, one may be able to solve a few more instances, but the largest grid sizes in the benchmark set remain out of reach for current exact methods.
@@ 4.5.1 Background
Figure 9: Oriented one-factorization of $K_{6}$ (left) and its associated single round-robin timetable (right), where each $i-j$ represents a home game of team $i$ against team $j$ for time slots $1$ to $5$.
@@ 4.5.2 Problem statement
Given a set of teams $T=\{1,\dots,n\}$ and time slots $S=\{1,\dots,2(n-1)\}$, $n$ even, a double round robin (2RR) corresponds to a perfect matching of teams in each time slot such that over all time slots each team plays precisely once at home against every other team. We call a 2RR ‘phased’ if each team meets every other team exactly once during the first and last $n-1$ time slots.
@@ 4.5.3 Instances
The problem instances stem from the International Timetabling Competition 2021 [226] and a follow-up study that generated additional problem instances using the same format [228]. These problem instances were generated using ‘Instance Space Analysis’ [213]. They feature a realistic set of constraints for real-world applications. Furthermore, they are challenging for state-of-the-art MIP and even dedicated sports tournament solvers. Finally, they feature a wide variety of characteristics that can expose different strengths and weaknesses of algorithms; see, e.g., [228]. Moreover, the instances were constructed so that a feasible solution is known to exist. More specifically, we made use of the fact that generating a basic sports schedule is easy, even when some of the constraints are considered. Suppose a problem instance should suffice $p$ ‘easy’ constraints from type A and $q$ ‘difficult’ constraints from type B. Then, we started by generating a solution that satisfies constraints A and added around this solution constraints of type B–such that the respective solution is compatible with constraints of type B.
The problem instances in this benchmark can be divided into four subclasses: ITC2021, Large, Medium, and Small. The ITC2021 class corresponds to the 45 original problem instances from the International Timetabling Competition on Sports Timetabling 2021 [118], considering all hard and soft constraints. Within the ITC2021 class are multiple instances for which no feasible solution was reported within a reasonable amount of time. The Large, Medium, and Small classes represent feasibility versions of the ITC2021 problem instances. More specifically, to lower the entrance barrier and simplify the problem, we ignore all soft constraints and, therefore, reduce the optimization problem to its feasibility version. For further experimentation, we also generated an additional set of problem instances having exactly the same number of constraints as the original ones, however, with 12 and 8 teams only. These instances are contained in Medium and Small.
@@ 4.5.4 Classical Baseline
Figure 10 shows the average solution times for a state-of-the-art Simulated Annealing (SA) algorithm (see [200]) and for a MIP formulation similar to the one presented in [26]. Both algorithms were run on a machine equipped with an AMD EPYC 7532 32-core processor, using 8 cores for solving the MIP (CPLEX 12.10) and a single core for the SA algorithm. The MIP solver was given a time limit of 24 hours, while the SA algorithm was stopped after a fixed number of iterations. The MIP performed reasonably well on small and most medium-sized instances but failed to solve any of the large-sized ones. Also, the SA algorithm managed to find only three large-sized instances, even though it was given a considerable amount of computational resources. As a reminder, the instances were selected in such a way that it is known that existing state-of-the-art solvers heavily struggle to find feasible solutions.
Figure 10: (Sports Scheduling) Solution times with Simulated Annealing (SA) and CPLEX (IP) for the feasibility version of the problem. The time limit for SA was iteration-based, while the IP solver was timed out after 24 hours.
@@ 4.6.1 Background
The problem of quantitative investment and risk diversification in portfolio optimization spans a broad spectrum of methodologies, each incorporating different simplifications, investor priorities, and trade-offs between risk exposure, returns, computational complexity, robustness, and out-of-sample performance. A foundational approach is Markowitz’s mean-variance optimization, formulated as a convex quadratic program [157]. This framework has been extended to address real-world constraints [138], estimation errors [98, 50, 141], sensitivity to input variability [172, 188, 35, 219], and Bayesian techniques that integrate expert judgment with statistical models [27, 12, 49]. Computational limitations of the methodologies for practically relevant cases [173] have led to the adoption of, e.g., fixed-mix strategies that improve model transparency and stability [166, 57], hierarchical risk parity and risk budgeting techniques [143, 94], and multiperiod approaches which help to capture intertemporal effects and hedging demands [195, 93]. In fact, portfolio optimization models can be defined with varying complexity [187, 136, 155, 108] depending on whether one aims to also jointly model risk, return predictability, and impact costs over time while addressing the exponential growth of time complexity with respect to the number of assets and rebalancing periods [33]. Portfolio optimization can transform into a stochastic dynamic programming problem, which, under realistic conditions that account for, e.g., transaction costs or short sell, has been proven to be computationally intractable for portfolios with more than two risky assets [45, 34].
@@
Given its relevance and complexity, portfolio optimization has attracted considerable interest from the quantum computing community. Theoretical research suggests that quantum algorithms may offer speedups under specific conditions that may, however, not hold in real-world instances; see, e.g., [135, 197, 152]. Another line of research has investigated how one may approach simple portfolio optimization instances with near-term compatible quantum algorithms [14, 214, 171]. The simplified version of the problem derived here focuses on the computational complexity in the multiperiod setting, using a binary quadratic formulation of the mean-variance model with linear constraints, and perfect knowledge of expected future returns and covariances. This allows for the creation of compact yet computationally challenging problem instances, and the systematic benchmarking of current quantum algorithmic capabilities.
@@
Why is it interesting? A vast array of industrial and scientific domains are interested in solving problems where one aims to find an optimal composition of a portfolio of coupled objects with respective observables that typically share a common probability space. (Financial) portfolio management is one of the hardest variants of this problem. It concerns the risk-sensitive allocation of a time-dependent investment budget for a number of diversified financial assets whose behavior varies with time according to unknown stochastic processes, are driven by external factors. The hardness of this problem can be explored at different levels of complexity, even at a small scale, and, as such, allows systematic tracking of algorithmic and hardware progress.
@@ 4.6.2 Problem statement
Let $\mathcal{T}:=\{t_{1},\dots,t_{m}\}$ be an ordered set of discrete time points such that $0<t_{1}<t_{2}<\cdots<t_{m}$, representing a finite investment horizon in the future. Consider an artificial deterministic market that is fully described by a one dimensional closed system of $n\in\mathbb{Z}^{+}$ financial assets $\{a_{i}\mid i=1\dots,n\}$, where each asset $a_{i}$ has known market prices $\{p_{it}\}$ at each time $t\in\mathcal{T}$. Assuming a myopic investor in a mean-variance framework and no access to additional information, the objective is to find a multiperiod trading strategy $\mathcal{S}_{0}:=\{x_{it}\}$ at $t_{0}$, where portfolio allocations $x_{it}\in\{0,1\}$ represent binary investment decisions of up to normalized capital units $C\in\mathbb{R}^{+}$ into assets $\{a_{i}\}$. The investor seeks to balance returns and risk at terminal time $t_{m}$, where the temporal returns are given by $(p_{it}-p_{i,t-1})x_{it}$ and the temporal risk is measured through the risk preference $\lambda$-weighted positive semidefinite covariance matrix $\Sigma$ as $\lambda(px)^{T}\Sigma xp$, while obeying a set of regularizing trading constraints.
@@
  • Transaction cost: Rebalancing of $x$ over $\mathcal{T}$ incurs costs due to fees, taxation, etc. that negatively offset the returns by a factor of transaction cost $\delta\in\mathbb{R}^{+}$ for each change $p_{it}\lvert x_{it-1}-x_{it}\rvert=p_{it}(x_{it-1}+x_{it}-2x_{it-1}x_{it})$ and a liquidation cost $\delta p_{i}x_{i}$ at terminal time $t_{m}$. This is a strong simplification since cost factors are neither global nor time-invariant scalars but rather stochastic functions driven by market uncertainties.
@@
  • Cash flow limits: Let $u\in\mathbb{R}^{+}$ denote the unit value of cash, and $C$ the total available units of normalized capital. Each asset is assigned an indicator $\tau_{i}\in\{-1,+1\}$, representing a short ($-1$) or long ($+1$) position. At any time, the total net investment $\sum_{i}\tau_{i}x_{it}$ must not exceed the capital limit $C$, and the total number of assets held at each time step must be bounded by $B$, which serves as an upper limit on the number of acquired positions.
@@
$\min_{{x\in\{0,1\}^{n\times t},y\in\{0,1\}^{c\times t}}}\quad$ $\sum_{t\in\mathcal{T}}\Bigg{(}\lambda\overbrace{\sum_{i}\sum_{j}p_{it}x_{it}\sigma_{ijt}x_{jt}p_{jt}}^{\text{risk}}$ (8)
$-\sum_{i}\Big{(}\underbrace{(p_{i,t+1}-p_{it})x_{it}}_{\text{return}}-\underbrace{\delta p_{it}(x_{it-1}+x_{it}-2x_{it-1}x_{it})}_{\text{transaction cost}}\Big{)}$
$\qquad-\underbrace{\nu u\sum_{c}2^{c}y_{ct}}_{\text{cash interest}}+\underbrace{\rho\sum_{f\in\mathcal{F}}p_{ft}x_{ft}}_{\text{short selling cost}}\Bigg{)}+\underbrace{\delta\sum_{i}p_{it_{m}}x_{it_{m}}}_{\text{liquidation cost}}$ (9)
@@
$\sum_{i}\tau_{i}x_{it}+\sum_{c}2^{c}y_{ct}$ $=C$ for all $\text{for all }t\in\mathcal{T}$ (10)
$\sum_{i}x_{it}$ $\leq B$ for all $\text{for all }t\in\mathcal{T}$ (11)
where (10) describes the capital limit and (11) limits the number of assets that can be included in the portfolio. Since we restricted the problem to binary variables, we modeled the amount of available cash units as binary variables $y_{ct}\in\{0,1\}$ with a suitable $c$ such that $\sum_{c}2^{(c+1)}>C$. It should be noted that other formulations may be possible as well.

### v2

@@ 4.4.2 Problem statement
For $n\in\mathbb{N}_{+}$, we write $n]=\{1,\ldots,n\}$. In the STPP, we are given an undirected simple graph $G=(V,E)$, and $n\in\mathbb{N}_{+}$ disjunct terminal sets $T_{1},\ldots,T_{n}$ with $T_{i}\subseteq V$ for all $i\in[n]$ and $T_{i}\cap T_{j}=\emptyset$ for $\in[n], iji\neq j$. The goal is to find an edge set $S^{*}\subseteq E$ of minimal cardinality such that $(V(S^{*}),S^{*})$ is a forest, where $\bigcup_{i\in[n]}T_{i}\subseteq V(S^{*})$ and no vertex from $T_{i}$ is connected to any vertex of $T_{j}$ for all $i\neq j$. The general STPP with $n>1$, as defined above, requires the inclusion of several disjoint Steiner trees into the same graph. A solution is a forest, and each terminal set $T_{i}$ is connected by a single tree in the forest.
@@ 4.4.3 Instances
We generated 3-dimensional $S\times S\times L$ grid graphs with holes. All terminals are placed at the border of the top layer. We require all trees to be node-disjoint. The parameters for each problem are the size of the grid graph $S$, the number of layers $L$, the maximum number of terminals per terminal set $T$, and the number of holes in the grid $H$. The holes can extend to multiple layers. An example of a $20\times 20\times 2$ instance with the corresponding optimal solution is given in Figure 7. In practice, the problem can also be defined with edge weights. For example, changing from one layer to another could be more expensive. However, since this does not make the problem more difficult in general, we kept it simple and set all edge weights to one. To ensure feasibility, the instances were reverse-engineered from feasible solutions, which may not be optimal. We observed that Gurobi started having difficulties with instances of size larger than $S=20$.
Finally, we provide a program that checks if a given solution to the STPP is feasible for a given instance, i.e., that each network’s terminals are connected by a Steiner Tree that is node-disjoint from all other Steiner Trees.
Figure 8: (Steiner) Total runtime and gap to the best-known (or even provably optimal) solution, solved with Gurobi for different size 2020 instances with a timeout of 72000s (left) and different size $30$ instances with a timeout of 86400s (right). The ticked line represents the timeout, the crosses indicate the runtime, and the stars show the gap.
@@ 4.4.4 Classical Baseline
Using the multicommodity flow ILP formulation for the node-disjoint STPP given in [110], we ran Gurobi 11.0.2 on an AMD EPYC-7513 32-core processor with 64 threads to attempt to solve the smallest instances in our benchmark set (gridsize 20 or 30). We used the model “as is”, without additional valid inequalities or decomposition strategies (the model is also provided in the repository [202]). Hardware information and results are summarized in Figure 8, whereas solution values and bounds are given in Table 4. With this ILP model, we could solve most of the size 20 instances (with an average solution time of 53 minutes) within 20 hours, but only two of the size 30 instances within 24 hours (the solution times were 194 minutes and 624 minutes). Reimplementing the techniques of [110] using current solver technology, one may be able to solve a few more instances, but the largest grid sizes in the benchmark set remain out of reach for current exact methods.
@@ 4.5.1 Background
Figure 9: Oriented one-factorization of $K_{6}$ (left) and its associated single round-robin timetable (right), where each iji-j represents a home game of team $i$ against team $j$ for time slots 11 to $5$.
@@ 4.5.2 Problem statement
Given a set of teams $T=\{1,\dots,n\}$ and time slots $S=\{1,\dots,2(n-1)\}$, $n$ even, a double round robin (2RR) corresponds to a perfect matching of teams in each time slot such that over all time slots each team plays precisely once at home against every other team. We call a 2RR ‘phased’ if each team meets every other team exactly once during the first and last n1n-1 time slots.
@@ 4.5.3 Instances
The problem instances stem from the International Timetabling Competition 2021 [235] and a follow-up study that generated additional problem instances using the same format [237]. These problem instances were generated using ‘Instance Space Analysis’ [222]. They feature a realistic set of constraints for real-world applications. Furthermore, they are challenging for state-of-the-art MIP and even dedicated sports tournament solvers. Finally, they feature a wide variety of characteristics that can expose different strengths and weaknesses of algorithms; see, e.g., [237]. Moreover, the instances were constructed so that a feasible solution is known to exist. More specifically, we made use of the fact that generating a basic sports schedule is easy, even when some of the constraints are considered. Suppose a problem instance should suffice $p$ ‘easy’ constraints from type A and $q$ ‘difficult’ constraints from type B. Then, we started by generating a solution that satisfies constraints A and added constraints of type B to this solution–such that the respective solution is compatible with constraints of type B.
The problem instances in this benchmark can be divided into four subclasses: ITC2021, Large, Medium, and Small. The ITC2021 class corresponds to the 45 original problem instances from the International Timetabling Competition on Sports Timetabling 2021 [124], considering all hard and soft constraints. Within the ITC2021 class are multiple instances for which no feasible solution was reported within a reasonable amount of time. The Large, Medium, and Small classes represent feasibility versions of the ITC2021 problem instances. More specifically, to lower the entrance barrier and simplify the problem, we ignore all soft constraints and, therefore, reduce the optimization problem to its feasibility version. For further experimentation, we also generated an additional set of problem instances having exactly the same number of constraints as the original ones; however, with 6, 8, and 12 teams only. These instances are contained in Medium and Small.
@@ 4.5.4 Classical Baseline
Figure 10 shows the average solution times for a state-of-the-art Simulated Annealing (SA) algorithm (see [209]) and for a MIP formulation similar to the one presented in [29]. Both algorithms were run on a machine equipped with an AMD EPYC-7532 32-core processor, using 8 cores for solving the MIP (CPLEX 12.10) and a single core for the SA algorithm. The MIP solver was given a time limit of 24 hours, while the SA algorithm was stopped after a fixed number of iterations. The MIP performed reasonably well on small and most medium-sized instances but failed to solve any of the large instances. Also, the SA algorithm managed to find only solutions to three large instances, even though it was given considerable computational resources. As a reminder, the instances were selected in such a way that it is known that existing state-of-the-art solvers heavily struggle to find feasible solutions.
Figure 10: (Sports Scheduling) Solution times with Simulated Annealing (SA) and CPLEX (ILP) for the feasibility version of the problem. The time limit for SA was iteration-based, while the ILP solver timed out after 24 hours.
@@ 4.6.1 Background
The problem of quantitative investment and risk diversification in portfolio optimization spans a broad spectrum of methodologies, each incorporating different simplifications, investor priorities, and trade-offs between risk exposure, returns, computational complexity, robustness, and out-of-sample performance. A foundational approach is Markowitz’s mean-variance optimization, formulated as a convex quadratic program [166]. This framework has been extended to address real-world constraints [145], estimation errors [105, 54, 149], sensitivity to input variability [181, 197, 39, 228], and Bayesian techniques that integrate expert judgment with statistical models [30, 12, 53]. Computational limitations of the methodologies for practically relevant cases [182] have led to the adoption of, e.g., fixed-mix strategies that improve model transparency and stability [175, 62], hierarchical risk parity and risk budgeting techniques [151, 101], and multiperiod approaches which help to capture intertemporal effects and hedging demands [204, 100]. In fact, portfolio optimization models can be defined with varying complexity [196, 143, 164, 115] depending on whether one aims to also jointly model risk, return predictability, and impact costs over time while addressing the exponential growth of time complexity with respect to the number of assets and rebalancing periods [37]. Portfolio optimization can transform into a stochastic dynamic programming problem, which, under realistic conditions that account for, e.g., transaction costs or short selling, has been proven to be computationally intractable for portfolios with more than two risky assets [49, 38].
@@
Given its relevance and complexity, portfolio optimization has attracted considerable interest from the quantum computing community. Theoretical research suggests that quantum algorithms may offer speedups under specific conditions that may, however, not hold in real-world instances; see, e.g., [142, 206, 162]. Another line of research has investigated how one may approach simple portfolio optimization instances with near-term compatible quantum algorithms [14, 223, 180]. The simplified version of the problem derived here focuses on the computational complexity in the multiperiod setting, using a binary quadratic formulation of the mean-variance model with linear constraints, and perfect knowledge of expected future returns and covariances. This allows for the creation of compact yet computationally challenging problem instances, as the problem is NP-hard [26], and the systematic benchmarking of current quantum algorithmic capabilities.
@@
Why is it interesting? A vast array of industrial and scientific domains are interested in solving problems where one aims to find an optimal composition of a portfolio of coupled objects with respective observables that typically share a common probability space. (Financial) Portfolio management is one of the hardest variants of this problem. It concerns the risk-sensitive allocation of a time-dependent investment budget for a number of diversified financial assets whose behavior varies with time according to unknown stochastic processes, which are driven by external factors. The hardness of this problem can be explored at different levels of complexity, even at a small scale, and, as such, allows systematic tracking of algorithmic and hardware progress.
@@ 4.6.2 Problem statement
Let $\mathcal{T}:=\{t_{1},\dots,t_{m}\}$ be an ordered set of discrete time points such that $0\leq t_{1}<t_{2}<\cdots<t_{m}$, representing a finite investment horizon in the future. Consider an artificial deterministic market that is fully described by a closed system of $n\in\mathbb{Z}^{+}$ financial assets $\{a_{i}\mid i=1\dots,n\}$, where each asset $a_{i}$ has known expected market prices $\{\bar{p}_{it}\}$ at each time $t\in\mathcal{T}$. Assuming a myopic investor in a mean-variance framework and no access to additional information, the objective is to find a multiperiod trading strategy $\mathcal{S}_{1}:=\{x_{it}\}$ at $t_{1}$, where portfolio allocations $x_{it}\in\{0,1\}$ represent binary investment decisions of up to $C$ normalized capital units $u\in\mathbb{R}^{+}$ into assets $\{a_{i}\}$. We define the price $p_{it}$ as follows: Let $e_{i}:=u/\bar{p}_{it_{0}}$, then $p_{it}=e_{i}\bar{p}_{it}$, i.e., we compute how many assets of type $i$ we can buy for one unit of cash at time $t_{1}$. The amount of stocks $e_{i}$ we can get at time $t_{1}$ for one unit of cash $u$ is then used for all later time steps, i.e., $p_{it}$ tells us what has become out of one cash unit invested into asset $i$ at time $t$. The investor seeks to balance returns and risk at terminal time $t_{m}$, where the temporal returns are given by $(p_{it}-p_{i,t-1})x_{it}$ and the temporal risk is measured through the risk preference $\lambda$-weighted positive semidefinite covariance matrix $\Sigma$ as $\lambda(px)^{T}\Sigma xp$, while obeying a set of regularizing trading constraints.
@@
  • Transaction cost: Rebalancing of $x$ over $\mathcal{T}$ incurs costs due to fees, taxation, etc. that negatively offset the returns by a factor of transaction cost $\delta\in\mathbb{R}^{+}$ for each change $p_{it}\lvert x_{it-1}-x_{it}\rvert=p_{it}(x_{it-1}+x_{it}-2x_{it-1}x_{it})$ and a liquidation cost $\delta p_{i}x_{i}$ at terminal time $t_{m}$. This is a strong simplification, as cost factors are neither global nor time-invariant scalars, but rather stochastic functions driven by market uncertainties.
@@
  • Cash flow limits: Let $u\in\mathbb{R}^{+}$ denote the unit value of cash, and $C$ the total available units of normalized capital. Each asset is assigned an indicator $\tau_{i}\in\{-1,+1\}$, representing a short (1-1) or long (+1+1) position. At any time, the total net investment $\sum_{i}\tau_{i}x_{it}$ must not exceed the capital limit $C$, and the total number of assets held at each time step must be bounded by $B$, which serves as an upper limit on the number of acquired positions.
@@
$\min_{{x\in\{0,1\}^{n\times m},y\in\{0,1\}^{|\mathcal{C}|\times m}}}\quad$ $\sum^{m}_{t=1}\Bigg{(}\lambda\overbrace{\sum^{n}_{i=1}\sum^{n}_{j=1}\tau_{i}p_{it}x_{it}\sigma_{ijt}\tau_{j}p_{jt}x_{jt}}^{\text{risk}}$ (8)
$-\sum^{n-1}_{i=1}\underbrace{\tau_{i}(p_{i,t+1}-p_{it})x_{it}}_{\text{return}}-\sum^{n}_{i=2}\underbrace{\delta p_{it}(x_{it-1}+x_{it}-2x_{it-1}x_{it})}_{\text{transaction cost}}$
$\qquad-\underbrace{\nu u\sum_{c\in\mathcal{C}}2^{c}y_{ct}}_{\text{cash interest}}+\underbrace{\rho\sum_{f\in\mathcal{F}}p_{ft}x_{ft}}_{\text{short selling cost}}\Bigg{)}+\underbrace{\delta\sum^{n}_{i=1}(p_{i1}x_{i1}+p_{im}x_{im})}_{t_{1}\text{ transaction $+$ liquidation cost}}$ (9)
@@
$\sum^{n}_{i=1}\tau_{i}x_{it}+\sum_{c\in\mathcal{C}}2^{c}y_{ct}$ $=C$ $\text{for all }t\in\mathcal{T}$ (10)
$\sum^{n}_{i=1}x_{it}$ $\leq B$ $\text{for all }t\in\mathcal{T}$ (11)
where (10) describes the capital limit and (11) limits the number of assets that can be included in the portfolio. Since we restricted the problem to binary variables, we modeled the amount of available cash units as binary variables $y_{ct}\in\{0,1\}$ with a suitable $c\in\mathcal{C}:=\{0,\ldots,\lfloor\log_{2}(C)\rfloor\}$ such that $\sum_{c\in\mathcal{C}}2^{c}\geq C$. It is worth noting that alternative formulations may also be possible.